OpenStudy (xofallenangelox):

Which property is shown? 3(–7 +12) = 3 • (–7) + 3 • (12) A. opposite of a sum property B. distributive property C. commutative property D. associative property

OpenStudy (xofallenangelox):

Will fan and medal

OpenStudy (anonymous):

Hey there! Do you know the definition of each of the answer choices?

OpenStudy (xofallenangelox):

No not really... it's confusing me a lot ;/

OpenStudy (xofallenangelox):

Explain?

OpenStudy (anonymous):

This is the commutative property: a + b = b + a

OpenStudy (anonymous):

Does the property they give you look like that?

OpenStudy (anonymous):

You are distributing the equation. The answer is still the same, but you change the order/arrangement. This is a prime example of the distributive property.

OpenStudy (xofallenangelox):

No it doesn't. This is a test I'm taking. I thought it was commutative property. But I'm still confused of why

OpenStudy (anonymous):

http://www.coolmath.com/prealgebra/06-properties This website might help

OpenStudy (xofallenangelox):

Okay, thx.

OpenStudy (anonymous):

Hey @Tegolbus Welcome to OpenStudy! You might want to read the Code of Conduct. The Code of Conduct just states the rules that OpenStudy users must follow. One of them is to help guide the asker, not just give them the answer. http://openstudy.com/code-of-conduct You're new here, so I won't hold it against you for accidentally breaking it :)

OpenStudy (xofallenangelox):

Wait, this is a commutative property. It is saying that a + b = b + a, something related. Is that all I had to do? I didn't have to solve the equation or anything?

OpenStudy (anonymous):

You don't need to solve the expression at all. Yes, commutative property is a + b = b + a 3(–7 +12) = 3 • (–7) + 3 • (12) is NOT commutative property, however. Also, keep in mind that associative property is a + (b+ c) = (a + b) + c And, finally, distributive property is a(b + c) = ab + ac So, is 3(–7 +12) = 3 • (–7) + 3 • (12) commutative, associative, or distributive property?

OpenStudy (xofallenangelox):

Wow! And, okay. Thank you! I think it's distributive?

OpenStudy (xofallenangelox):

Lol, this is much easier, now. I'm going to say Distributive.

OpenStudy (anonymous):

Correct!

OpenStudy (xofallenangelox):

Yay!